# How does My Model Fail? Automatic Identification and Interpretation of Physical Plausibility Failure Modes with Matryoshka Transcoders

### v1

@@ 1 Introduction
Figure 1: Matryoshka Transcoders for Automatic Physical Plausibility Analysis in Generative Models. Our framework automatically discovers and interprets physical plausibility violations in generated images. Images from generative models pass through a physical plausibility classifier, whose intermediate representations are analyzed by Matryoshka Transcoders to extract hierarchical sparse features. These features are automatically interpreted by large multimodal models, producing natural language descriptions of diverse error patterns including anatomical impossibilities, spatial inconsistencies, gravity violations, and other physical implausibilities.
@@
Despite their prevalence and importance, physical plausibility errors remain challenging to detect systematically. Existing evaluation metrics primarily focus on semantic alignment between images and text [radford2021learningtransferablevisualmodels, sampaio2024typescoretextfidelitymetric, yao2024hifi] or aggregate distribution quality [yu2021frechet, jayasumana2024rethinkingfidbetterevaluation], leaving specific physical violations largely undetected [bianchi2024clipmainroadblockfinegrained, schrodi2025effectstriggermodalitygap]. Recent research has shown that even state-of-the-art large multimodal models (LMMs) struggle to consistently identify subtle physical implausibilities when evaluating images holistically [tang2025humanlikecontentanalysisgenerative]. This detection gap prevents developers from understanding where and how their models fail, hindering targeted improvements.
@@
Furthermore, no framework exists for automatically identifying and interpreting specific physical error patterns in natural language. While recent work has explored physical plausibility assessment through sparse coding [tang2025humanlikecontentanalysisgenerative], existing approaches suffer from limited feature accuracy and struggle to discover fine-grained error patterns. The absence of interpretable, automatically discovered features means that practitioners lack systematic understanding of the diverse ways generative models violate physical constraints, limiting the development of targeted mitigation strategies.
@@
We introduce Matryoshka Transcoders (See Figure 1), a novel framework for the automatic discovery and interpretation of physical plausibility features in generative models. Our approach extends the Matryoshka representation learning paradigm [bussmann2025learningmultilevelfeaturesmatryoshka] to transcoder architectures [dunefsky2024transcodersinterpretablellmfeature], enabling hierarchical sparse feature learning at multiple granularity levels, from coarse-grained violations to fine-grained specific errors. We train Matryoshka Transcoders on intermediate representations from a physical plausibility classifier trained on human-annotated datasets, then leverage LMMs to automatically interpret discovered features through a two-stage prompting strategy that first identifies common visual patterns and then analyzes their physical plausibility.
@@
Our experiments demonstrate that Matryoshka Transcoders successfully discover diverse physical plausibility patterns without manual feature engineering, achieving superior feature relevance and accuracy compared to existing methods including standard transcoders, sparse autoencoders, and prior interpretability approaches. We utilize these discovered patterns to establish the first interpretation-based benchmark for evaluating physical plausibility in generative models. Our analysis of eight state-of-the-art generative models, including SDXL variants [meng2022sdeditguidedimagesynthesis], DALL-E3 [dalle3], FLUX [FLUX.1-dev], Kolors [Kolors-diffusers], and Stable-Cascade [stablecascade2025], reveals systematic failure modes and provides actionable insights into how these models violate physical constraints, from textual distortions to anatomical impossibilities. These findings pave the way for targeted model improvements and more physically plausible generative systems.
@@
  • Feature Relevance Metrics. We introduce two complementary metrics to quantitatively evaluate how well discovered features align with human-assigned topics.
@@ 3.2 Physical Plausibility Classifier Training
We employ human annotators to systematically label a dataset of natural images according to their physical plausibility. Each image is assigned a binary label indicating whether it contains physical plausibility errors or is physically plausible. To ensure annotation quality and consistency, we provide annotators with clear guidelines defining physical plausibility violations and include representative examples of common error types. The annotated dataset is augmented with images from natural image datasets, including MSCOCO [lin2015microsoftcococommonobjects] and Flickr8k [plummer2016flickr30kentitiescollectingregiontophrase], which serve as additional negative samples of physically plausible images.
@@ 3.4 LMM Interpretation
Stage 2: Physical plausibility analysis. Given the identified summarization, we then prompt the LMM to determine whether this pattern represents a physical plausibility violation and, if so, to describe the nature of the violation in natural language.
@@ 4.2 Physical Plausibility Pattern Identification
Figure 3 presents representative discovered features with their automatically generated interpretations. Our method identifies diverse violation patterns in varied error types: Feature 104 captures anatomical distortions in tennis scenes, Feature 243 detects malformed text in posters, and Feature 648 recognizes implausible hand-drawn styles. Each feature demonstrates semantic coherence across its activating images, with natural language descriptions accurately characterizing the specific physical plausibility violation.
@@
We evaluate feature quality using two complementary relevance metrics that assess both the selectivity of discovered features and the quality of their natural language interpretations.
Population-based relevance. For each feature $i$ with activation threshold $\theta$, we compute its topic-relevance score as:
@@
Table 1 shows that Matryoshka Transcoders outperform all baselines on both relevance metrics, achieving 16.00% population-based and 17.52% description-based scores. This demonstrates the effectiveness of combining targeted transcoder learning with hierarchical Matryoshka training for discovering selective physical plausibility features.
@@
SAE 6.54% 39.96%
Matryoshka SAE 9.88% 41.71%
Transcoder 12.74% 34.70%
Matryoshka Transcoder 16.00% 42.07%
@@
Figure 4: Feature relevance distribution across methods. Percentage distribution of features by wrong ratio across four sparse dictionary learning methods. Features above the relevance threshold of 0.95 are considered as task-relevant features.
@@ 4.3 Generative Model Benchmark
Model Name Mean Error Feature Count
@@
Table 2: Physical plausibility evaluation of state-of-the-art generative models. Error count represents the average number of physical plausibility features activated per generated image across 5,000 test samples. Lower values indicate better performances. FLUX achieves the best physical plausibility, while SDXL-turbo exhibits substantially more violations than other models.

### v2

@@ 1 Introduction
Figure 1: Matryoshka Transcoders for Automatic Physical Plausibility Analysis in Generative Models. We pass images from generative models to a physical plausibility classifier, whose representations are analyzed by Matryoshka Transcoders to extract hierarchical sparse features related to physical plausibility. These features encompasses anatomical impossibilities, spatial inconsistencies, gravity violations, and other physical implausibilities and are interpreted by LMMs automatically.
@@
Despite their prevalence and importance, physical plausibility errors remain challenging to detect systematically. Existing evaluation metrics primarily focus on semantic alignment between images and text [radford2021learningtransferablevisualmodels, sampaio2024typescoretextfidelitymetric, yao2024hifi] or aggregate distribution quality [yu2021frechet, jayasumana2024rethinkingfidbetterevaluation], leaving specific physical violations largely undetected [bianchi2024clipmainroadblockfinegrained, schrodi2025effectstriggermodalitygap]. Recent research has shown that even state-of-the-art large multimodal models (LMMs) struggle to consistently identify subtle physical implausibilities when evaluating images holistically [tang2025humanlikecontentanalysisgenerative], motivating us to develop methods that can identify such failure modes of generative models. This detection gap leaves such failure modes largely unidentified, hindering the understanding of how some generated images fail in physical plausibility.
@@
Furthermore, no framework exists for automatically identifying and interpreting specific physical error patterns in natural language. Though recent work [tang2025humanlikecontentanalysisgenerative] has proposed to explore physical plausibility detection by the identification of topic-related visual patterns grounded in natural language, they utilize preliminary interpretability modules and struggle to discover fine-grained error patterns. The absence of interpretable, automatically discovered features means that practitioners lack systematic understanding of the diverse ways generative models violate physical constraints, limiting the development of targeted mitigation strategies.
@@
We introduce Matryoshka Transcoders (See Figure 1), a novel framework for the automatic discovery and interpretation of physical plausibility features in generative models. Our approach extends the Matryoshka representation learning paradigm [bussmann2025learningmultilevelfeaturesmatryoshka] to transcoder architectures [dunefsky2024transcodersinterpretablellmfeature], enabling hierarchical sparse feature learning at multiple granularity levels, from coarse-grained violations to fine-grained specific errors. We first assign human annotators to annotate a physical plausibility dataset, including a set of generated images that are marked as physically plausible or not. We use this dataset to train a physical plausibility classifier that contains physical plausibility features, and develop Matryoshka Transcoders on the latent representations of the classifier. The Matryoshka Transcoder identifies specific physical plausibility features, which are then interpreted by LMMs to get their human understandable explanations.
@@
Our experiments demonstrate that Matryoshka Transcoders successfully discover diverse physical plausibility patterns without manual feature engineering, achieving superior feature relevance and accuracy compared to existing methods including standard transcoders, sparse autoencoders, and prior interpretability approaches. We utilize these identified patterns related to physical plausibility to establish the first interpretation-based benchmark for evaluating physical plausibility in generative models. Our analysis of eight state-of-the-art generative models, including SDXL variants [meng2022sdeditguidedimagesynthesis], DALL-E3 [dalle3], FLUX [FLUX.1-dev], Kolors [Kolors-diffusers], and Stable-Cascade [stablecascade2025], reveals systematic failure modes and provides actionable insights into how these models violate physical constraints, from textual distortions to anatomical impossibilities. These findings pave the way for targeted model improvements and more physically plausible generative systems.
@@
  • Feature Relevance Metrics. We introduce two metrics, population-based relevance and description-based relevance, to quantitatively evaluate how well discovered features align with human-assigned topics.
@@ 3.2 Physical Plausibility Classifier Training
We employ two independent human annotators to label a dataset of generated images according to their human understanding on the physical plausibility of these images. Each image is assigned a binary label indicating whether it contains physical plausibility errors or is physically plausible. To ensure annotation quality and consistency, we provide annotators with clear guidelines defining physical plausibility violations and include representative examples of common error types. The annotated dataset is augmented with images from natural image datasets, including MSCOCO [lin2015microsoftcococommonobjects] and Flickr8k [plummer2016flickr30kentitiescollectingregiontophrase], which serve as additional negative samples of physically plausible images.
@@ 3.4 LMM Interpretation
Stage 2: Physical plausibility analysis. Given the identified summarization, we prompt the LMM to determine whether this pattern represents a physical plausibility violation and, if so, to describe the violation in natural language.
@@ 4.2 Physical Plausibility Pattern Identification
Figure 3 presents representative discovered features with their automatically generated interpretations. Our method identifies diverse violation patterns in varied error types: Feature 104 captures anatomical distortions in tennis scenes, Feature 243 detects malformed text in posters, and Feature 648 recognizes implausible hand-drawn styles. Each feature demonstrates semantic coherence across its activating images, with natural language descriptions characterizing the specific physical plausibility violation.
@@
We evaluate feature quality using two complementary relevance metrics that assess both the selectivity of discovered features and the relevance of their description.
Population-based relevance. For each feature $i$ with activation threshold $\theta$, we compute its topic-relevance score (wrong ratio) as:
@@
We benchmark our method against sparse autoencoders [makhzani2014ksparseautoencoders], transcoders [dunefsky2024transcodersinterpretablellmfeature], and Matryoshka SAEs [bussmann2025learningmultilevelfeaturesmatryoshka] as baselines. Specifically, we utilize a large corpus of images generated by different models with diverse captions to identify samples that activate each feature in these interpretability modules, enabling us to compute their relevance scores. Table 1 shows that Matryoshka Transcoders outperform all baselines on both relevance metrics, achieving 16.00% population-based and 17.52% description-based scores. This demonstrates the effectiveness of combining targeted transcoder learning with hierarchical Matryoshka training for discovering physical plausibility features.
@@
SAE 6.54% 6.01%
Matryoshka SAE 9.88% 5.66%
Transcoder 12.74% 11.52%
Matryoshka Transcoder 16.00% 13.04%
@@
Figure 4: Feature relevance distribution across methods. Percentage distribution of features by their opic-relevance score (wrong ratio) across four sparse dictionary learning methods. Features above the relevance threshold of 0.95 are considered as task-relevant features.
@@ 4.3 Generative Model Benchmark
For each generative model, we compute the mean error feature count as:
$\text{Error Count}=\frac{1}{N}\sum_{j=1}^{N}|\{i:f_{i}(x_{j})>\theta\land M_{i}\geq 0.95\}|$ (7)
where $N$ is the number of generated images, and $M_{i}$ is the topic-relevance score. This metric counts how many topic-relevant features activate on each image, averaged across all images. Lower scores indicate better performance, as fewer error-selective features respond to the model’s outputs.
Model Name Mean Error Count
@@
Table 2: Physical plausibility evaluation of state-of-the-art generative models. Error count represents the average number of physical plausibility features activated per generated image across 5,000 test samples. Lower values indicate better performances. FLUX achieves the best physical plausibility, while SDXL-turbo exhibits substantially more violations than other models. Image samples of these generative models can be found in Appendix.
@@ 7 Conclusion
A. Qualitative Examples of the Generative Model We Evaluate
To complement our quantitative evaluation in Table 2, we present qualitative samples from all eight evaluated generative models in Figure 6. For each model, we selected a diverse set of representative images spanning common generation scenarios. These examples validate that our metric captures meaningful differences in physical plausibility that align with human perception of image quality.
Figure 6: Qualitative samples from evaluated generative models ranked by physical plausibility. Each row shows representative outputs from a different model, ordered by mean error count. Lower error counts correspond to better visual quality and fewer physical violations, with FLUX achieving the best performance (38.0) and SDXL-turbo exhibiting the most plausibility failures (74.2).
B. Full LMM Prompts
In this section, we fully describe the prompts we use in our pipelines in detail.
Pattern Summarization Prompt ($\text{Prompt}_{\text{sum}}$) You are an expert in multimodal feature analysis. Given the following images and their captions: {image-text pairs}. Analyze the commonalities among these images. Identify: If there exist one common feature that is possessed by all the instances. Summarize and output exactly one feature, for example: ’[Commonality: Animal wildlife in natural habitats]’ and ’[Commonality: Strawberry-based dessert or dish]’. Only answer in general, do not analyze each image one by one, only generate one single concise phrase. Start answer with ’[Commonality:’. End with ’]’
Error Interpretation Prompt ($\text{Prompt}_{\text{interp}}$) You are an expert in analyzing visual content for physical plausibility errors, anatomical accuracy, and generation artifacts in AI-generated images. I will show youlen(samples)len(samples) sample images from a learned feature in an interpretability module trained to detect physical plausibility errors. **Feature Commonality**: ”commonalitycommonality” **Error Ratio**: $error_{ratio}$} of these images contain physical errors ($error_{count}/total_{count}$). Be strict with the evaluations. Usually only error ratios larger than 0.95 indicate clear physical plausibility failure mode as these features should not be activated in the presence of any correct images. Other features can be unusable and should be considered as no such error mode. Here are the sample images: {image-text pairs}. Based on the $len(samples)$ images shown above and knowing that: - Feature commonality: ”commonality” - $error_{ratio}$} contain physical plausibility errors Analyze whether these images share a **common physical plausibility error or anatomical inaccuracy**. Focus ONLY on: **Physical Errors:** - Incorrect number of fingers, toes, or limbs - Extra or missing body parts - Distorted anatomy or impossible body proportions - Unnatural poses or joint configurations - Incorrect object physics (floating, defying gravity) - Impossible spatial arrangements or perspectives - Anatomically incorrect faces or features - Object inconsistencies or impossible constructions **Important:** - Ignore style, artistic choices, or image quality - Ignore semantic content unless it relates to physical errors - Focus ONLY on violations of physical or anatomical plausibility **Response Format:** If there IS a common physical error across these images: [Error: Brief description of the specific error] If there are NO clear physical plausibility errors for all images or if the error ratio does not suggest clear and monosemantic error pattern: [No common errors] Your response MUST start with either ”[Error:” or ”[No common errors]” and end with ”]”.
C. Effect of Dictionary Size on Feature Distribution.
Figure 7 demonstrates how dictionary size affects the quality and quantity of discovered physical plausibility features. We trained Matryoshka Transcoders with five nested dictionary sizes ranging from 1,024 to 16,384 latents and analyzed the distribution of features by their topic-relevance scores. Smaller dictionaries (1,024 latents) achieve higher feature relevance scores (17.2%), producing more concentrated and reliable features, but discover fewer total relevant patterns (170 features). In contrast, larger dictionaries (16,384 latents) provide broader coverage of error types, discovering 1,896 relevant features, but with lower average relevance (11.6%). Notably, all configurations exhibit a consistent bimodal distribution with features clustering at 70-80% and above 95% relevance, revealing a natural bottleneck that separates task-correlated features from truly topic-specific features. The average wrong ratio remains stable (67-70%) across all dictionary sizes, and nearly all features remain active (94-98%), indicating efficient dictionary utilization. This analysis motivates our nested training approach with $\mathcal{M}=\{128,256,512,1024,2048\}$, which combines the high relevance of small dictionaries with the comprehensive coverage of larger ones.
Figure 7: Feature relevance distribution across dictionary sizes in Matryoshka Transcoders. Distribution of features by their topic-relevance score (wrong ratio) for five nested dictionary sizes (1024, 4096, 8192, 16384 latents). Features above the relevance threshold of $0.95$ (indicated by dashed line) are considered topic-relevant. Smaller dictionaries yield higher concentration of relevant features ($17.2\%$ for 1024) but discover fewer total patterns, while larger dictionaries achieve broader coverage (1,896 relevant features for 16384) with lower relevance scores ($11.6\%$). The bimodal distribution consistently shows feature clustering at $70-80\%$ and $\geq 95\%$ relevance ranges across all configurations, revealing a natural bottleneck between task-correlated and topic-specific features.